# Monomial arrow removal and the finitistic dimension conjecture

### v1

@@ Abstract.
Finitistic dimension, Bound quiver algebras, Monomial arrow removal, Cleft extension of abelian categories, Gröbner bases
@@ 2.1. Monomial arrow removal
We begin by showing that the only possible paths $p$ and $q$ such that there is a single $\alpha$-monomial generating set $T=T_{p,q}$ of $I$ are $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$ or $q=\beta\gamma_{2}\delta_{2}$. The maximal non-zero paths of $\Lambda$ that end at vertex $1$ are the paths $\lambda_{1}^{j_{1}-1}$, $\lambda_{9}^{j_{9}-1}\theta$ and $\zeta\eta\theta$. Therefore, path $p$ must divide one of the above paths from the right. Similarly, the maximal non-zero paths of $\Lambda_{1}$ starting at vertex 2 are the paths $\lambda_{2}^{j_{2}-1}$, $\beta\gamma_{1}\delta_{1}\lambda_{6}^{j_{6}-1}$ and $\beta\gamma_{1}\delta_{1}\varepsilon$, and $q$ has to divide one of them, or their variants where the subpath $\gamma_{1}\delta_{1}$ is replaced by the path $\gamma_{2}\delta_{2}$, from the left. If $p$ divides $\lambda_{1}^{j_{1}-1}$ and $q$ is any of the aforementioned candidate paths, then $p\alpha q\in I$ implies that $q\in I$, a contradiction. Therefore, path $p$ is divided by $\theta$ from the right. A similar argument shows that $q$ cannot be a subpath of $\lambda_{2}^{j_{2}-1}$. Furthermore, if we take $q$ to be $\beta\gamma_{1}$ or $\beta\gamma_{2}$ then, for any candidate path $p$, it holds that $p\alpha q\notin I$. We deduce that $q$ has to be divided by $\beta\gamma_{1}\delta_{1}$ or $\beta\gamma_{2}\delta_{2}$ from the left. Since $\theta\alpha\beta\gamma_{1}\delta_{1}=\theta\alpha\beta\gamma_{2}\delta_{2}\in I$, it holds that $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$ or $q=\beta\gamma_{2}\delta_{2}$.
Note that the above arguments show that $T=T_{p,q}$ is a single $\alpha$-monomial generating set for $I$, where $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$. Indeed, it follows that no proper subpath of $p\alpha q$ is in $I$. However, the generating set $T$ is not strict $\alpha$-monomial as the path $\gamma_{1}\delta_{1}$ overlaps from the left with and divides $q$. Moreover, the ideal $I$ possesses no strict $\alpha$-monomial generating set as the paths $\gamma_{1}\delta_{1}$ and $\gamma_{2}\delta_{2}$ occur in every generating set for $I_{1}$, their length being $2$. In particular, the quotient algebra $\Lambda/\langle\overline{\alpha}\rangle$ is a single but not strict monomial arrow removal of $\Lambda$.
@@
We close this subsection by showing that, if $\Lambda/\langle\overline{\alpha}\rangle$ is a monomial arrow removal of $\Lambda$, then the two algebras form a cleft extension. The following result, whose proof is included for the sake of completeness, is a particular case of [6, Lemma 3.1].
@@ 4. Rule IV:
Assume now that $z,w\in kQ$ and there are paths $u$, $v$ such that $\operatorname{\mathrm{Tip}}(z)v=u\operatorname{\mathrm{Tip}}(w)$ and $\operatorname{\mathrm{Tip}}(z)$, $\operatorname{\mathrm{Tip}}(w)$ do not divide $u$ and $v$, respectively. In other words, we assume that $\operatorname{\mathrm{Tip}}(z)$ overlaps with $\operatorname{\mathrm{Tip}}(w)$ from the right. Then the overlapping relation of $(z,w)$ by $(u,v)$ is
@@
    the set $\operatorname{\mathrm{Tip}}(\mathcal{G})$ satisfies properties (P2), (P2)${}^{\textrm{op}}$ and (P3).
@@
Let now $\mathcal{G}_{\textrm{red}}$ be the reduced Gröbner basis of $I$, and observe that $\operatorname{\mathrm{Tip}}(\mathcal{G}_{\textrm{red}})$ satisfies properties (P2), (P2)${}^{\textrm{op}}$ and (P3), as $\operatorname{\mathrm{Tip}}(\mathcal{G}_{\textrm{red}})\subseteq\operatorname{\mathrm{Tip}}(\mathcal{G})$. It remains to show that $p\alpha q\in\mathcal{G}_{\textrm{red}}$ and $\mathcal{G}_{\textrm{red}}\setminus\{p\alpha q\}$ avoids $\alpha$.
@@
(P2)${}^{\textrm{op}}$ If $q$ overlaps with $sv$ from the right, then either $q$ overlaps with $s$ from the right or $s$ divides $q$. This is a contradiction in either case, as $X$ satisfies properties (P2)${}^{\textrm{op}}$ and (P3).
@@
(P2) If $p$ overlaps with $h$ from the left, then $p$ overlaps with $csd$ from the left for some path $s$ occurring in $f$. There are three cases. If $p$ overlaps with $d$ from the left, then $p$ overlaps with $\operatorname{\mathrm{Tip}}(g)$ from the left. If $p$ does not overlap with $d$ from the left but it does so with $sd$, then $p$ overlaps with $s$ from the left. If $p$ does not overlap with $d$ or $sd$ from the left then it does so with $csd$, in which case $s$ divides $p$. All three cases contradict the assumption that the set $\{f,g\}$ satisfies properties (P2) and (P3). We omit the proof of property (P2)${}^{\textrm{op}}$ as it is analogous to the proof of property (P2).
@@
Recall that $T=T_{p,q}$ is the reduced Gröbner basis for $I$ with respect to any admissible order on $\mathcal{B}_{Q}$ such that $\gamma_{1}\delta_{1}$ is smaller than $\gamma_{2}\delta_{2}$. It is easy to see that $T=T_{p,q}$ is strict $\alpha$-monomial as a Gröbner basis by verifying property (P1) for $T$ and properties (P2), (P2)${}^{\textrm{op}}$ and (P3) for $\operatorname{\mathrm{Tip}}(T)$. It should be noted that $T$ is not strict $\alpha$-monomial as a generating set for $I$.
@@ 5. Examples
    all simple modules have infinite injective dimension and projective dimension at least $2$;
@@
First, all simple modules have infinite projective and injective dimension due to the loops $\lambda_{i}$ according to the Strong No Loop Theorem [10]. Furthermore, the algebra $\Lambda$ is triangle reduced as the quiver $Q$ is strongly connected, see [6, Lemma 3.12]. Lastly, every arrow of $Q$ occurs in every generating set $T$ for $I$. Indeed, every arrow different from $\alpha$ occurs in $I$ in a path of length $2$ and, therefore, it must occur in $T$ as well. On the other hand, if $\alpha$ did not occur in $T$, then $\langle\overline{\alpha}\rangle$ would be a projective $\Lambda$-bimodule according to [9], and this would contradict the fact that $I$ possesses a strict $\alpha$-monomial Gröbner basis. To see that, note that the homomorphism of $\Lambda$-bimodules $\Lambda{e_{1}}\otimes_{k}{e_{2}}\Lambda\to\langle\overline{\alpha}\rangle$ such that ${e_{1}}\otimes{e_{2}}\mapsto\overline{\alpha}$ is a projective cover, and its kernel contains the non-zero element $\overline{p}\otimes\overline{q}$.
We also note that $\operatorname{\mathsf{fin.dim}}\Lambda\neq 0$ according to [2, Lemma 6.2], since the left regular module ${}_{\Lambda}\Lambda$ does not possess a submodule isomorphic to the simple left module corresponding to vertex $2$. Indeed, no path occurring in $T$ starts with $\alpha$, and thus $z\in I$ whenever $z\in{e_{2}{kQ}}$ and $\alpha z\in I$. Furthermore, the algebra $\Lambda$ is not monomial, and its Loewy length can be arbitrarily large as integers $j_{i}\geq 2$ may be chosen freely. Finally, if we choose $j_{7}=2$ (i.e. $\lambda_{7}^{2}\in I$), then [6, Corollary 4.12] applied to the loop $\lambda_{7}$ and the arrow $\varepsilon$ implies that $\Lambda$ is not Iwanaga-Gorenstein, since the paths $\varepsilon\lambda_{7}$ and $\varepsilon\zeta$ are zero in $\Lambda$.
@@
through an application of 4.5. Indeed, 4.5 implies directly that $\operatorname{\mathsf{fin.dim}}\Lambda\leq\operatorname{\mathsf{fin.dim}}{\Lambda/\langle\overline{\alpha}\rangle}+2$. Furthermore, we perform seven consecutive triangle reductions for the quotient algebra $\Lambda/{\langle\overline{\alpha}\rangle}$ (see [5]). At every step, we extract a local algebra corresponding to the external vertex we want to remove (the unique source or sink vertex), or the direct product of two local algebras corresponding to vertices $4$ and $5$, until we reach two local algebras corresponding to the remaining two vertices. As each such reduction can increase the finitistic dimension by at most one, and local algebras have zero finitistic dimension, this process implies that $\operatorname{\mathsf{fin.dim}}{\Lambda/\langle\overline{\alpha}\rangle}\leq 7$. Consequently, the upper bound $\operatorname{\mathsf{fin.dim}}\Lambda\leq 9$ holds.
@@
First, we argue that all simple modules have projective dimension at least $2$. Let $S_{i}$ denote the simple module associated to vertex $i$ for $i=1,2,\ldots,6$. Computations then show that
$\operatorname{\mathsf{pd}}_{\Lambda}S_{2}$ $=4,$
$\operatorname{\mathsf{pd}}_{\Lambda}S_{6}$ $=2,$
@@
In addition to the above, all arrows of $Q$ occur in every generating set for $I$, and no triangular reduction is possible as the quiver $Q$ is strongly connected, see [6, Lemma 3.12]. Moreover, it holds that $\Lambda$ is not monomial as the paths $\gamma\varepsilon_{1}$ and $\gamma\varepsilon_{2}$ are equal and non-zero, the Loewy length of $\Lambda$ is $8$ where $\theta_{2}\alpha\varepsilon_{1}\zeta\theta_{1}\alpha\varepsilon_{2}$ is the longest non-zero path, and $\operatorname{\mathsf{fin.dim}}\Lambda$ is non-zero as the simple left $\Lambda$-module corresponding to vertex $3$, for instance, is not isomorphic to a submodule of ${}_{\Lambda}\Lambda$ ([2, Lemma 6.2]).
We show next that the algebra $\Lambda$ admits the quotient algebra $\Lambda/\langle\overline{\alpha}\rangle$ as a strict monomial arrow removal. Specifically, the generating set $T=T_{p,q}$ is strict $\alpha$-monomial for $p=\theta_{1}$ and $q=\varepsilon_{1}$ according to 2.6. Indeed, property (P1) clearly holds for $T$ and $\theta_{1}\alpha\varepsilon_{1}\in T$. Property (P3) also holds as no path occurring in $T$ may divide $p$ or $q$, being the case that they comprise a single arrow each. Furthermore, properties (P2) and (P2)${}^{\textrm{op}}$ hold as no path occurring in $T$ starts with arrow $\varepsilon_{1}$ or ends with arrow $\theta_{1}$.
@@
where $S_{9}$ is the simple right $\Lambda$-module corresponding to vertex $9$. But $\operatorname{\mathsf{pd}}S_{9}=\infty$ due to the Strong No Loop Theorem [10] and our claim follows. In particular, 4.5 does not apply for $\Lambda$ and the arrow $\alpha$.

### v2

@@ Abstract.
Finitistic dimension, Bound quiver algebras, Monomial arrow removal, Cleft extensions, (non-commutative) Gröbner bases
@@ 1. Introduction and main result
Acknowledgments
The present research project was partly supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.). Specifically, the second author received support under the “3rd Call for H.F.R.I. Ph.D. Fellowships” (FN: 47510/03.04.2022). For the third author, the research project was implemented in the framework of the H.F.R.I. call “Basic Research Financing (Horizontal Support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union - NextGenerationEU (H.F.R.I. Project Number: 16785).
@@ 2.1. Monomial arrow removal
We begin by showing that the only possible paths $p$ and $q$ such that there is a single $\alpha$-monomial generating set $T=T_{p,q}$ of $I$ are $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$ or $q=\beta\gamma_{2}\delta_{2}$. The maximal non-zero paths of $\Lambda$ that end at vertex 111 are the paths $\lambda_{1}^{j_{1}-1}$, $\lambda_{9}^{j_{9}-1}\theta$ and $\zeta\eta\theta$. Therefore, path $p$ must divide one of the above paths from the right. Similarly, the maximal non-zero paths of $\Lambda_{1}$ starting at vertex 2 are the paths $\lambda_{2}^{j_{2}-1}$, $\beta\gamma_{1}\delta_{1}\lambda_{6}^{j_{6}-1}$ and $\beta\gamma_{1}\delta_{1}\varepsilon$, and $q$ has to divide one of them, or their variants where the subpath $\gamma_{1}\delta_{1}$ is replaced by the path $\gamma_{2}\delta_{2}$, from the left. If $p$ divides $\lambda_{1}^{j_{1}-1}$ and $q$ is any of the aforementioned candidate paths, then $p\alpha q\in I$ implies that $q\in I$, a contradiction. Therefore, path $p$ is divided by $\theta$ from the right. A similar argument shows that $q$ cannot be a subpath of $\lambda_{2}^{j_{2}-1}$. Furthermore, if we take $q$ to be $\beta\gamma_{1}$ or $\beta\gamma_{2}$ then, for any candidate path $p$, it holds that $p\alpha q\notin I$. We deduce that $q$ has to be divided by $\beta\gamma_{1}\delta_{1}$ or $\beta\gamma_{2}\delta_{2}$ from the left. Since $\theta\alpha\beta\gamma_{1}\delta_{1}=\theta\alpha\beta\gamma_{2}\delta_{2}\in I$, it holds that $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$ or $q=\beta\gamma_{2}\delta_{2}$.
Note that the above arguments show that $T=T_{p,q}$ is a single $\alpha$-monomial generating set for $I$, where $p=\theta$ and $q=\beta\gamma_{1}\delta_{1}$. Indeed, it follows that no proper subpath of $p\alpha q$ is in $I$. However, the generating set $T$ is not strict $\alpha$-monomial as the path $\gamma_{1}\delta_{1}$ overlaps from the left with and divides $q$. Moreover, the ideal $I$ possesses no strict $\alpha$-monomial generating set as the paths $\gamma_{1}\delta_{1}$ and $\gamma_{2}\delta_{2}$ occur in every generating set for $I_{1}$, their length being 222. In particular, the quotient algebra $\Lambda/\langle\overline{\alpha}\rangle$ is a single but not strict monomial arrow removal of $\Lambda$.
@@
We close this subsection by showing that, if $\Lambda/\langle\overline{\alpha}\rangle$ is a monomial arrow removal of $\Lambda$, then the two algebras form a cleft extension. The following result, whose proof is included for the sake of completeness, is a particular case of various results found in [6, Sections 3 and 4].
@@ 4. Rule IV:
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5. Examples
    all simple modules have infinite injective dimension and projective dimension at least 222;
@@
First, all simple modules have infinite projective and injective dimension due to the loops $\lambda_{i}$ according to the Strong No Loop Theorem [10]. Furthermore, the algebra $\Lambda$ is triangle reduced as the quiver $Q$ is strongly connected, see [6, Lemma 3.12]. Lastly, every arrow of $Q$ occurs in every generating set $T$ for $I$. Indeed, every arrow different from $\alpha$ occurs in $I$ in a path of length 222 and, therefore, it must occur in $T$ as well. On the other hand, if $\alpha$ did not occur in $T$, then $\langle\overline{\alpha}\rangle$ would be a projective $\Lambda$-bimodule according to [9], and this would contradict the fact that $I$ possesses a strict $\alpha$-monomial Gröbner basis. To see that, note that the homomorphism of $\Lambda$-bimodules $\Lambda{e_{1}}\otimes_{k}{e_{2}}\Lambda\to\langle\overline{\alpha}\rangle$ such that ${e_{1}}\otimes{e_{2}}\mapsto\overline{\alpha}$ is a projective cover, and its kernel contains the non-zero element $\overline{p}\otimes\overline{q}$.
We also note that $\operatorname{\mathsf{fin.dim}}\Lambda\neq 0$ according to [2, Lemma 6.2], since the left regular module ${}_{\Lambda}\Lambda$ does not possess a submodule isomorphic to the simple left module corresponding to vertex 222. Indeed, no path occurring in $T$ starts with $\alpha$, and thus $z\in I$ whenever $z\in{e_{2}{kQ}}$ and $\alpha z\in I$. Furthermore, the algebra $\Lambda$ is not monomial, and its Loewy length can be arbitrarily large as integers $j_{i}\geq 2$ may be chosen freely. Finally, if we choose $j_{7}=2$ (i.e. $\lambda_{7}^{2}\in I$), then [6, Corollary 4.14] applied to the loop $\lambda_{7}$ and the arrow $\varepsilon$ implies that $\Lambda$ is not Iwanaga-Gorenstein, since the paths $\varepsilon\lambda_{7}$ and $\varepsilon\zeta$ are zero in $\Lambda$.
@@
through an application of 4.5. Indeed, 4.5 implies directly that $\operatorname{\mathsf{fin.dim}}\Lambda\leq\operatorname{\mathsf{fin.dim}}{\Lambda/\langle\overline{\alpha}\rangle}+2$. Furthermore, we perform seven consecutive triangle reductions for the quotient algebra $\Lambda/{\langle\overline{\alpha}\rangle}$ (see [5]). At every step, we extract a local algebra corresponding to the external vertex we want to remove (the unique source or sink vertex), or the direct product of two local algebras corresponding to vertices 444 and 555, until we reach two local algebras corresponding to the remaining two vertices. As each such reduction can increase the finitistic dimension by at most one, and local algebras have zero finitistic dimension, this process implies that $\operatorname{\mathsf{fin.dim}}{\Lambda/\langle\overline{\alpha}\rangle}\leq 7$. Consequently, the upper bound $\operatorname{\mathsf{fin.dim}}\Lambda\leq 9$ holds.
@@
First, we argue that all simple modules have projective dimension at least 222. Let $S_{i}$ denote the simple module associated to vertex $i$ for $i=1,2,\ldots,6$. Computations then show that
$\operatorname{\mathsf{pd}}_{\Lambda}S_{2}$ =4=4,= 4 ,
$\operatorname{\mathsf{pd}}_{\Lambda}S_{6}$ =2=2,= 2 ,
@@
In addition to the above, all arrows of $Q$ occur in every generating set for $I$, and no triangular reduction is possible as the quiver $Q$ is strongly connected, see [6, Lemma 3.12]. Moreover, it holds that $\Lambda$ is not monomial as the paths $\gamma\varepsilon_{1}$ and $\gamma\varepsilon_{2}$ are equal and non-zero, the Loewy length of $\Lambda$ is 888 where $\theta_{2}\alpha\varepsilon_{1}\zeta\theta_{1}\alpha\varepsilon_{2}$ is the longest non-zero path, and $\operatorname{\mathsf{fin.dim}}\Lambda$ is non-zero as the simple left $\Lambda$-module corresponding to vertex 333, for instance, is not isomorphic to a submodule of ${}_{\Lambda}\Lambda$ ([2, Lemma 6.2]).
[…]
@@
where $S_{9}$ is the simple right $\Lambda$-module corresponding to vertex 999. But $\operatorname{\mathsf{pd}}S_{9}=\infty$ due to the Strong No Loop Theorem [10] and our claim follows. In particular, 4.5 does not apply for $\Lambda$ and the arrow $\alpha$.